OpenStudy (anonymous):

The table below shows some input and output values for a function named F(n). Which of the following equations describes F(n)? n F(n) 1 9 2 19 5 49 8 79 10 99 20 199 A. F(n) = n + 8 B. F(n) = 10n - 1 C. F(n) = 10n + 9

OpenStudy (danjs):

It looks like each increase of 1 for n, results in an increase of 10 for F(n) and F(1) = 9 from the table F(n) = 10n - 1 Is the one that works with both those requirements

OpenStudy (danjs):

f(1) = 10(1) - 1 = 9 f(2) = 10*2 - 1 = 19 . . .

OpenStudy (danjs):

each output, it one less than 10 times the input